MANCOSA: POSTGRADUATE DIPLOMA IN BUSINESS MANAGEMENT 59 Additional information New shares were issued on 1 January 2018.

DOLL 58 MANCOSA: POSTGRADUATE DIPLOMA IN BUSINESS MANAGEMENT DUE DATE: 26 AUGUST 2019 ASSESSMENT 5 CCOUNTING AND FINANCIAL MANAGEMENT (20) QUESTION1 The following information was extracted from the accounting records Statement of Comprehensive Income for the year ended 31 December 2017 (R) 2018 (R) 1 856 000 1 2000 Sales (750 000 (1 280 000) F3 Cost of sales 450 000 576 000 Gross profit (212 000 (291 200) Operating expenses 26 30000 Depreciation 2 186 0 261 200 Other operating expenses 238 0 284 800 Operating profit (72 000 QW (24 000) Interest on mortgage loan 166 260 800 Profit before tax (58 100 90 240) Income tax S A 107 9 170 560 Profit after tax Stątement of Financial Position as at 31 December 2017 (R 2018 (R ASSETS 2 444 2 464 160 Non-current assets roeeement 2 444 00 2 464 1 336 3560 Current assets 250 64 00 120 nventories (all Trading Inventory Trade & other receivables 22 0 196 0 Cash & cash equivalents 2 780 00 2 820 1601 2017 (R 2018 (R EQUITY AND LIABILITIES Equity 1 980 0 2 437 5 1 760 00 2 180 0 Ordinary share capital (issued at R10 each) 22000 257 56 Retained income 600 000 200 Non-current liability (Mortgage loan, 12% pa) Current liabilities 200 00 182 60 138 000 98 00 Trade & other payables South African Revenue Services (income tax) Shareholders for dividends 6 00 8 6 56 000 76 00 2 780 2 820 160 58 PROGRAMME HANDBOOK: JULY 2018 INTAKE
MANCOSA: POSTGRADUATE DIPLOMA IN BUSINESS MANAGEMENT 59 Additional information " New shares were issued on 1 January 2018. " Interim and final dividends for the year ended 31 December 2018 amounted to R133 000 s Fixed assets were sold at carrying (book) value during the year for R144 000. p PgtD Study the information provided above by Belle Mare Ltd and prepare the Statement of Cash Flows (using the indirect method) for the year ended 31 December 2018. (20) QUESTION 2 The Statement of Comprehensive Income provided below was obtained from the accounting records of Statement of Comprehensive Income for the year ended 31 August 2018 (Extract) 8 5 000 Sales 5 Cost of sales (1 900000 Gross profit 3 100 (700 000) Operating expenses 2 Operating profit 2 400 Interest income 50 Interest expense Profit before tax 2 200 (630 000) Income tax Net profit 1 570 Required: 2.1 Explain SIX (6) reasons why shareholders and financial managers would be interested in the Statement of Comprehensive Income. Calculate the interest expense for the year ended 31 August 2018. 2.2 Explain the accounting treatment of sales returns and allowances. 2.3 2.4 Interest has been calculated at 8% per year on the amount invested. Assume that there was no change in investments during the year, determine the value of the investment. Which method of depreciation will result in lower net profit in the early years of the life of the asset? Explain why 2.5 Of what significance is the operating profit to investors? 2.6 59 PROGRAMME HANDBOOK: JULY 2018 INTAKE
DOLL COSA: POSTGRADUATE DIPLOMA İN BUSINESS MANAGEMENT 60 QUESTION 3 universal Ltd plans to manufacture a new product and the following information is applicable (20) Estimated sales for the year 2019 40 000 units at R140 each Estimated costs for the year 2019 Direct material R72 per unit Direct labour R12 per unit Factory overheads (all fixed) R110 000 per annum Selling expenses 30% of sales Administrative expenses (all fixed) R180 000 per annum Required Calculate the following independently: 3.1 Break-even quantity 3.2 Break-even value. 3.3 Break-even quantity, if the direct labour cost is increased by R2 per unit 3.4 Selling price per unit, if the profit per unit is R12. 3.5 New break-even quantity and value if the selling price is increased by 5%. 5 (20) QUESTION 4 2 Traders is considering selling a refrigerator to a customer on credit. The cost of the refrigerator is the selling price is R10 000. A credit term of 2/10 net 30 was agreed upon. The cost of capital to Rover Traders is 15%. Required: Use the information given above to answer the following questions: 4.1.1 Calculate the profit that Rover Traders would make if the account is settled within 10 days. 4.1.2 Should the customer fail to pay the amount due and the account is written off after 90 days, what would be the loss to the firm? The annual sales of product Y of Govender Limited is 400 000 units. The purchase price is R12 per unit. The carrying costofproductYamountst 30%ofthe unit purchase rice. The ordering cost is R45 per order Required Use the information provided above by Govender Limited to calculate the: 4.2.1 Economic order quantity (EOQ) 4.2.2 Number of orders that need to be placed each year. 4.3 or one year. The stated interest rate is 15% per annum ansto borrow R1 000 000 f Jumbo Enterprises Required Use the information provided above to calculate the effective interest rate if 4.3.1 the interest is discounted 4.3.2 there is a 25% compensating balance requirement GEFO 60 PROGRAMME HANDBOOK: JULY 2018 INTAKE
MANCOSA: POSTGRADUATE DIPLOMA IN BUSINESS MANAGEMENT 41 ION S Enterprises has the option to invest in machinery in Projects M and N but finance is only vabeo 2os one of them Project M Project N IR) Initial cost 450 450 000 Net profit: 36 000 69 000 Year 1 Year 2 75 000 69 000 Year 3 102 000 69 000 Year 4 129 000 69 000 1. Assume that all cash flows take place at the end of the year except the original investment in the project which takes place at the beginning of the project. 2. Project M machinery is expected to be disposed of at the end of year 5 with a scrap value of R60 000 3. Project N machinery is expected to be disposed of at the end of year 5 with a nil scrap value. 4. Depreciation is calculated on a straight-line basis. 5. The discount rate to be used by the company is 12%. 5.1 Required: Use the information provided above by Dolby Enterprises to answer the following questions: 5.1.1 Calculate the Payback Period of Project N. (Answer must be expressed in years and months.) 5.1.2 Calculate the Accounting Rate of Return (on average investment) of Project M. (Answer must be expressed to two decimal places.) 5.1.3 Calculate the Net Present Value of each project. (Round off amounts to the nearest Rand.) 5.1.4 Using your answers from question 5.1.3, which project should be chosen? Why? 5.2 A machine with a purchase price of R418 000 is estimated to eliminate manual operations and save the company R130 000 cash per year. The machine will last 5 years and have no residual value at the end of its useful life Required Calculate the Internal Rate of Return (answer expressed to two decimal places).
Answer #1

Solution to Question No: ;3

Tha data is given as follows:

Sales = 40,000 units

Selling Price(SP) =  R140 each

Sales Value(SV) =(40,000x140)= 56,00,000

Variable costs(VC):

Direct material = 72 per unit

Direct labor= 12 per unit

Selling expenses= (140x30%)= 42 per unit

VC per unit =  126 per unit

Total Variable cost= (40000x126)=50,40,000

Contribution = 56,00,000-50,40,000= 560,000

Fixed cost (FC):

Factory overheads = 110,000

Administration exps = 180,000

Total FC = 290,000

3.1: Break even Quantity:

= (FC/Contribution per unit)

Contribution per unit = SP-VC = 140-126 =14

=(290,000/14) = 20714.28

= 20714 units

3.2: Break Even Value:

= (FC x SV)/(Contribution)

=(290,000 x 56,00,000)/560,000

=29,00,000

OR

= Break even units X Selling price per unit

= 20714 x 140 = 28,99,960 = 29,00,000

3.3: Break even quantity, if the direct labor cost increased by R2 per unit :

So the new variable cost will be = 126+2 =128 per unit

New BEP quantity:

   = (FC/Contribution per unit)

Contribution per unit = SP-VC = 140-128 =12

=(290,000/12) = 24166.66

= 24167 units

3.4: Selling price per unit, if the profit per unit R12

Profit = Contribution per unit - FC per unit

(Fixed cost per unit =Total FC/Units= 290,000/40000=7.25)

12 = Contribution per unit - 7.25

Contribution per unit= 12+7.25 = 19.25

We know that Original Contibution margin is 10% (Contribution/SP =14/140=10%)

So new selling price will be = 19.25 x 10 = 192.5

3.5: New Break even quantity and value , if the selling price increased by 5%

New Selling price = 140+5% = 147 per unit

New Variable costs(VC):

Direct material = 72 per unit

Direct labor= 12 per unit

Selling expenses= (147x30%)= 44.1 per unit

VC per unit = 128.1 per unit

New Contribution = 147 - 128.1 = 18.9

Break even Quantity:

= (FC/Contribution per unit)

=(290,000/18.9) = 15343.91

= 15344 units

Break Even Value:

= Break even units X Selling price per unit

= 15344 x 147 = 22,55,568